OpenStudy (anonymous):

Points of inflection for y=tan^-1(x)

OpenStudy (anonymous):

\[dy/dx=1/(1+x^2)\] This can never equal zero because x^2 can never be less than 0. \[d^2y/dx^2=2x/(1+x^2)^2\]. This equals zero when x=0. so there is a point of inflection at x=0.